# Spectral gap bounds for reversible hybrid Gibbs chains

### v1

@@ Abstract
Hybrid Gibbs samplers represent a prominent class of approximated Gibbs algorithms that utilize Markov chains to approximate conditional distributions, with the Metropolis-within-Gibbs algorithm standing out as a well-known example. Despite their widespread use in both statistical and non-statistical applications, very little is known about their convergence properties. This article introduces novel methods for establishing bounds on the convergence rates of hybrid Gibbs samplers. In particular, we examine the convergence characteristics of hybrid random-scan Gibbs and data augmentation algorithms. Our analysis reveals that the absolute spectral gap of a reversible hybrid chain can be bounded based on the absolute spectral gap of the exact Gibbs chain and the absolute spectral gaps of the Markov chains employed for conditional distribution approximations. The new techniques are applied to four algorithms: a random-scan Metropolis-within-Gibbs sampler, a hybrid proximal sampler, random-scan Gibbs samplers with block updates, and a hybrid slice sampler.
@@ 1 Introduction
Although these hybrid algorithms have enjoyed pervasive popularity and extensive application, there remains a notable scarcity of analytical techniques tailored for understanding their convergence properties. This article aims to address this gap. Intuitively, if the Markovian steps in a hybrid Gibbs sampler approximate the intractable conditional distributions well, then the convergence rate of the hybrid Gibbs sampler should be similar to that of the corresponding exact Gibbs sampler. Our analysis confirms this intuition and makes this notion precise with some quantitative convergence rate inequalities. In particular, we provide bounds on the absolute spectral gap of the hybrid algorithm in terms of that of the exact algorithm and the quality of the approximation. Bounds involving Dirichlet forms are also provided. We focus on the case where both the exact Gibbs and the hybrid algorithms are reversible. To be precise, we study random-scan Gibbs and data augmentation algorithms. Moreover, the Markovian approximation of a conditional distribution is assumed to be reversible. The efficiency of a reversible MCMC sampler is closely related to its spectral gap. For a Markov chain associated with a Markov transition kernel (Mtk) $K$ that is reversible with respect to a distribution $\omega$, its absolute spectral gap is defined to be $1-\|K\|_{\omega}$, where $\|K\|_{\omega}\in[0,1]$ is the $L^{2}$ norm of $K$ when $K$ is viewed as a Markov operator. (See Section 3 for the exact definitions.) The $L^{2}$ distance between the distribution of the chain’s $t$th iteration and its target $\omega$ decreases at a rate of $\|K\|_{\omega}^{t}$ as $t$ tends to infinity. Indeed, the norm $\|K\|_{\omega}$ can be regarded as the convergence rate of the chain, with a smaller rate indicating faster convergence. The number of iterations it takes for the chain to reach within some $L^{2}$ distance of its stationary distribution is roughly proportional to $-1/\log\|K\|_{\omega}$, which is approximately $1/(1-\|K\|_{\omega})$ when $\|K\|_{\omega}$ is not too close to 0. Our results give quantitative upper and lower bounds on $(1-\|K_{2}\|_{\omega})/(1-\|K_{1}\|_{\omega})$ when $K_{1}$ and $K_{2}$ are the Mtks of a Gibbs sampler and its hybrid version. In Section 5, we apply the spectral gap bounds, derived under general conditions, to four examples, illustrating their relevance and potential for various applications. Together, these examples effectively showcase how our framework enriches and diversifies the analysis of exact and hybrid Gibbs samplers across different scenarios.
@@
  • Firstly, our method can be efficiently integrated with existing techniques to establish new and explicit convergence bounds. In Section 5.1, we study a random-scan Gibbs sampler whose conditional distributions are approximated using a random-walk Metropolis algorithm. We derive an explicit relationship between the spectral gaps of the exact and hybrid samplers when the conditional distributions are well-conditioned using recent studies on log-concave sampling (Andrieu et al., 2022b). In Section 5.2, we study a new type of proximal sampler (Lee et al., 2021a), which can be regarded as a data augmentation algorithm, or a deterministic-scan Gibbs algorithm with two components. We quantitatively bound its convergence rate when one of the conditional distributions involved is approximated via a random-walk Metropolis step.
@@
We now review some relevant existing works. On finite state spaces, Liu (1996) constructed a Metropolized random-scan Gibbs sampler and showed that it is asymptotically more efficient than the corresponding exact Gibbs sampler using Peskun’s ordering. Analysis on general state spaces is more challenging. The convergence properties of hybrid Gibbs samplers were studied qualitatively by Roberts and Rosenthal (1997), Fort et al. (2003), and Qin and Jones (2022). There have been very few quantitative studies on the subject. When the approximating Markov chains are uniformly ergodic with convergence rates bounded away from one, some quantitative convergence bounds are given in Roberts and Rosenthal (1998), Neath and Jones (2009), and Jones et al. (2014). The bounds herein rely on the weaker assumption that the approximating chains are geometrically ergodic with convergence rates bounded away from unity. In fact, one of our bounds for hybrid data augmentation chains relies on yet a weaker condition. Some of the linear algebraic techniques that we employ are adopted from Qin and Wang (2023) and Łatuszyński and Rudolf (2014). The rest of this article is organized as follows. An overview of our results is given in Section 2. Section 3 gives a brief survey on the $L^{2}$ theory of Markov chains, and contains the definitions of concepts like Markov operator and spectral gap. The main results are formally stated in Section 4. In Section 5 we apply our main result to four examples. Section 6 contains a brief discussion on avenues for future research. The Appendix contains some technical results and proofs.
@@ 2.1 Gibbs and hybrid Gibbs samplers
We begin by formally defining conditional distributions through the notion of disintegration. Let $(\mathsf{X},\mathcal{B},\Pi)$ be a probability space. For $i=1,\dots,n$ where $n$ is a positive integer, let $(\mathsf{Y}_{i},\mathcal{F}_{i})$ be a measurable space, and let $\psi_{i}:\mathsf{X}\to\mathsf{Y}_{i}$ be a measurable function. Let $M_{i}=\Pi\circ\psi_{i}^{-1}$, and let $(P_{i,y_{i}})_{y_{i}\in\mathsf{Y}_{i}}$ be a disintegration of $\Pi$ with respect to $\psi_{i}$. That is, (i) for $y_{i}\in\mathsf{Y}_{i}$, $P_{i,y_{i}}(\cdot)$ is a $\sigma$-finite measure, and, for $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, $P_{i,y_{i}}(\cdot)$ is a probability measure concentrated on the pre-image $\psi^{-1}_{i}(y_{i}):=\psi^{-1}_{i}(\{y_{i}\})=\{x\in\mathsf{X}:\,\psi_{i}(x)=y_{i}\}$, (ii) for $f:\mathsf{X}\to[0,\infty)$ measurable, $y_{i}\mapsto\int_{\mathsf{X}}f(x)\,P_{i,y_{i}}(\mathrm{d}x)$ is measurable, and
@@ 2.2 A loose summary of main results
We now give an overview of our main results. The exact statements are given in Section 4, after a brief survey on the $L^{2}$ theory of Markov chains in Section 3. For now, recall that the $L^{2}$ norm of a Markov operator is in $[0,1]$. When the associated chain is reversible, the norm can be viewed as its convergence rate, with a smaller rate indicating faster convergence to the stationary distribution. It is also common to quantify the convergence speed of chains through the spectral gap, which is one minus the operator norm. The norms of $T$, $\hat{T}$, $S$, $\hat{S}$, and $Q_{i,y_{i}}$ are denoted by $\|T\|_{\Pi}$, $\|\hat{T}\|_{\Pi}$, $\|S\|_{\Pi}$, $\|\hat{S}\|_{\Pi}$, and $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}$, respectively. For the random-scan Gibbs sampler, our bound loosely goes like this:
@@
$\left(\inf_{i}\text{Spectral gap of chain }C_{i}\right)\times\text{Spectral gap of chain }A\leq\text{Spectral gap of chain }B.$
Interestingly, bounds of this form have appeared in various other settings that are quite distinct from the one at hand (Carlen et al., 2003; Madras and Randall, 2002; Qin, 2023). Previously, the only quantitative convergence bound for general hybrid random-scan Gibbs samplers that we know of is from Roberts and Rosenthal (1998). In the context of Example 3, Roberts and Rosenthal (1998) established a relation that bears some similarity to the first inequality in (2). However, their proof relied on the strong assumptions that the chain associated with $Q_{i,y_{i}}$ is uniformly ergodic for all $i$ and $y_{i}$, and the chain associated with $T$ is uniformly ergodic as well.
For the data augmentation sampler, we have two sets of bounds. The first is reminiscent of (2). Loosely speaking, it goes like this:
$\left(1-\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\right)(1-\|S\|_{\Pi})\leq 1-\|\hat{S}\|_{\Pi}\leq\left(1+\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{i,y_{i}}}\right)(1-\|S\|_{\Pi}).$ (3)
If the Markov operator $Q_{2,y_{2}}$ is positive semi-definite for each $y_{2}$, then the right-hand-side can be tightened to $1-\|S\|_{\Pi}$. The first inequality in (3) generalizes Theorem 6 of Jones et al. (2014). The proof of that theorem leads to a bound that is similar to the first inequality in (3), but it holds only if the chain associated with $Q_{2,y_{2}}$ is uniformly ergodic for every $y_{2}$. The second set of bounds loosely goes like this: for any positive even integer $t$,
$t(1-\|\hat{S}\|_{\Pi})\geq 1-\|\hat{S}\|_{\Pi}^{t}\geq 1-\|S\|_{\Pi}-\sup_{y_{1}}\int_{\mathsf{X}}\|Q_{2,\psi_{2}(x)}\|_{P_{2,\psi_{2}(x)}}^{t}P_{1,y_{1}}(\mathrm{d}x).$ (4)
If $Q_{2,y_{2}}$ is positive semi-definite for each $y_{2}$, then (4) hold for odd values of $t$ as well. The first inequality in (4) is a version of Bernoulli’s inequality. When $\|S\|_{\Pi}<1$, the second inequality is nontrivial when
@@
This is a weaker condition than $\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}<1$, which must hold for the lower bound on $1-\|\hat{S}\|_{\Pi}$ in (3) to be non-trivial. However, in terms of bounding $1-\|\hat{S}\|_{\Pi}$ from below, (3) can be sharper than (4) if $\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}<1$. The lower bound on $1-\|\hat{S}\|_{\Pi}$ in (4) is a generalization of a result in Łatuszyński and Rudolf (2014), who studied a hybrid slice sampler. In Section 5.4, it is shown that this hybrid slice sampler is in fact a special case of the hybrid data augmentation sampler.
@@
We were not able to obtain an analogue of (4) for the random-scan Gibbs chain that was sufficiently simple and useful.
@@ 3 Preliminaries
This section reviews the $L^{2}$ theory for Markov chains. Let $(\mathsf{Y},\mathcal{F},\omega)$ be a probability space. The linear space $L^{2}(\omega)$ consists of functions square-integrable function $f:\mathsf{Y}\to\mathbb{R}$ such that $\int_{\mathsf{Y}}f(y)^{2}\,\omega(\mathrm{d}y)<\infty$. Two functions in $L^{2}(\omega)$ are considered identical if they are equal almost everywhere. For $f,g\in L^{2}(\omega)$ and $c\in\mathbb{R}$, $(f+g)(y)=f(y)+g(y)$, and $(cf)(y)=cf(y)$. The inner product of $f$ and $g$ is $\langle f,g\rangle_{\omega}=\int_{\mathsf{Y}}f(y)g(y)\,\omega(\mathrm{d}y)$. The norm of $f$ is $\|f\|_{\omega}=\sqrt{\langle f,f\rangle_{\omega}}$. Let $L_{0}^{2}(\omega)$ be the subspace of $L^{2}(\omega)$ consisting of functions $f$ such that $\omega f:=\int_{\mathsf{Y}}f(y)\,\omega(\mathrm{d}y)=0$. It can be shown that $(L^{2}(\omega),\langle\cdot,\cdot,\rangle_{\omega})$ and $(L_{0}^{2}(\omega),\langle\cdot,\cdot,\rangle_{\omega})$ are real Hilbert spaces. A probability measure $\mu:\mathcal{F}\to[0,1]$ is said to be in $L_{*}^{2}(\omega)$ if $\mathrm{d}\mu/\mathrm{d}\omega$ exists and is in $L^{2}(\omega)$. The $L^{2}$ distance between $\mu,\nu\in L_{*}^{2}(\omega)$ is
@@
where $\mu f$ and $\nu f$ are defined analogously to $\omega f$. Let $K:\mathsf{Y}\times\mathcal{F}\to[0,1]$ be an Mtk such that
@@
We refer to $K$ as a Markov operator. The operator corresponding to the $t$-step Mtk $K^{t}(\cdot,\cdot)$ is precisely the operator $K^{t}=\prod_{i=1}^{t}K$. The operator $K$ is self-adjoint, i.e., $\langle Kf,g\rangle_{\omega}=\langle f,Kg\rangle_{\omega}$ for $f,g\in L_{0}^{2}(\omega)$, if and only if the corresponding chain is reversible with respect to $\omega$. We say $K$ is positive semi-definite if it is self-adjoint, and $\langle f,Kf\rangle_{\omega}\geq 0$ for $f\in L_{0}^{2}(\omega)$. With a bit of work, one can show that $T$ and $S$ from Section 2 are both positive semi-definite linear operators. By Jensen’s inequality,

### v2

@@ Abstract
Hybrid Gibbs samplers represent a prominent class of approximated Gibbs algorithms that utilize Markov chains to approximate conditional distributions, with the Metropolis-within-Gibbs algorithm standing out as a well-known example. Despite their widespread use in both statistical and non-statistical applications, relatively little is known about their convergence properties. This article gives novel bounds on the convergence rates of hybrid Gibbs samplers. In particular, we examine the convergence characteristics of hybrid random-scan Gibbs and data augmentation algorithms. Our analysis confirms that the absolute spectral gap of a hybrid chain can be bounded based on the absolute spectral gap of the exact Gibbs chain and the absolute spectral gaps of the Markov chains employed for conditional distribution approximations. For application, we study the convergence properties of four practical hybrid Gibbs algorithms: a random-scan Metropolis-within-Gibbs sampler, a hybrid proximal sampler, random-scan Gibbs samplers with block updates, and a hybrid slice sampler.
@@ 1 Introduction
Despite the pervasive popularity and extensive application of these hybrid algorithms, significant gaps persist in our understanding of their convergence properties. Following the pioneering works of Roberts and Rosenthal (1997) and Andrieu et al. (2018), this article endeavors to bridge the gap by establishing a connection between the convergence rates of exact and hybrid Gibbs chains in some important scenarios. Intuitively, if the Markovian steps in a hybrid Gibbs sampler approximate the intractable conditional distributions well, then the convergence rate of the hybrid Gibbs sampler should be similar to that of the corresponding exact Gibbs sampler. Our analysis confirms this intuition and makes this notion precise with some quantitative convergence rate inequalities. In particular, we provide bounds on the absolute spectral gap of the hybrid algorithm in terms of that of the exact algorithm and the quality of the approximation. Bounds involving Dirichlet forms are also provided.
We focus on the case where both the exact Gibbs and the hybrid algorithms are reversible. To be precise, we study random-scan Gibbs and data augmentation algorithms. Moreover, the Markovian approximation of a conditional distribution is assumed to be reversible.
@@
The efficiency of a reversible MCMC sampler is closely related to its spectral gap. For a Markov chain associated with a Markov transition kernel (Mtk) $K$ that is reversible with respect to a distribution $\omega$, its absolute spectral gap is defined to be $1-\|K\|_{\omega}$, where $\|K\|_{\omega}\in[0,1]$ is the $L^{2}$ norm of $K$ when $K$ is viewed as a Markov operator. (See Section 3 for the exact definitions.) The $L^{2}$ distance between the distribution of the chain’s $t$th iteration and its target $\omega$ decreases at a rate of $\|K\|_{\omega}^{t}$ as $t$ tends to infinity. Indeed, the norm $\|K\|_{\omega}$ can be regarded as the convergence rate of the chain, with a smaller rate indicating faster convergence. The number of iterations it takes for the chain to reach within some $L^{2}$ distance of its stationary distribution is roughly proportional to $-1/\log\|K\|_{\omega}$, which is approximately $1/(1-\|K\|_{\omega})$ when $\|K\|_{\omega}$ is not too close to 0. Our results give quantitative upper and lower bounds on $(1-\|K_{2}\|_{\omega})/(1-\|K_{1}\|_{\omega})$ when $K_{1}$ and $K_{2}$ are the Mtks of a Gibbs sampler and its hybrid version.
In Section 5, we apply the spectral gap bounds, derived under general conditions, to four examples, illustrating their relevance and potential for various applications. Together, these examples effectively showcase how our framework enriches and diversifies the analysis of exact and hybrid Gibbs samplers across different scenarios.
@@
  • Firstly, the methods herein can be efficiently integrated with existing techniques to establish new and explicit convergence bounds. In Section 5.1, we study a random-scan Gibbs sampler whose conditional distributions are approximated using a random-walk Metropolis algorithm. We derive an explicit relationship between the spectral gaps of the exact and hybrid samplers when the conditional distributions are well-conditioned using recent studies on log-concave sampling (Andrieu et al., 2022b). In Section 5.2, we study a new type of proximal sampler (Lee et al., 2021a), which can be regarded as a data augmentation algorithm, or a deterministic-scan Gibbs algorithm with two components. We quantitatively bound its convergence rate when one of the conditional distributions involved is approximated via a random-walk Metropolis step.
@@
We now review some relevant existing works and explain how our contribution relates to them. On finite state spaces, Liu (1996) constructed a Metropolized random-scan Gibbs sampler and showed that it is asymptotically more efficient than the corresponding exact Gibbs sampler using Peskun’s ordering. Analysis on general state spaces is more challenging. The convergence properties of hybrid Gibbs samplers were studied qualitatively by Roberts and Rosenthal (1997), Fort et al. (2003), and Qin and Jones (2022). Quantitative studies on the subject remain scarce. When the approximating Markov chains are uniformly ergodic with convergence rates bounded away from one, some quantitative convergence bounds are given in Roberts and Rosenthal (1998), Neath and Jones (2009), and Jones et al. (2014). For hybrid random-scan Gibbs algorithms, we develop convergence bounds under the weaker assumption that the approximating chains are geometrically ergodic with convergence rates bounded away from unity. This is our first main contribution. We also develop convergence bounds for hybrid data augmentation algorithms. For these algorithms, Andrieu et al. (2018) gave results leading to convergence bounds that are useful under conditions akin to those we use for our bounds for hybrid random-scan Gibbs algorithms. As our second main contribution, we develop convergence bounds for hybrid data augmentation algorithms that remain useful under even weaker conditions. Some of the linear algebraic techniques that we employ are adopted from Qin and Wang (2023) and Łatuszyński and Rudolf (2014). Following the initial release of our work, Ascolani et al. (2024) independently derived bounds for the $s$-conductance of a class of hybrid random-scan Gibbs algorithms, exhibiting some similarities to our spectral gap bounds.
The rest of this article is organized as follows. An overview of our results is given in Section 2. Section 3 gives a brief survey on the $L^{2}$ theory of Markov chains, and contains the definitions of concepts like Markov operator and spectral gap. The main results are formally stated in Section 4. In Section 5 we apply these results in Section 4 to four examples. Section 6 contains a brief discussion on avenues for future research. The Appendix contains some technical results and proofs.
@@ 2.1 Gibbs and hybrid Gibbs samplers
We begin by formally defining conditional distributions through the notion of disintegration. Let $(\mathsf{X},\mathcal{B},\Pi)$ be a probability space. For $i=1,\dots,n$ where $n$ is a positive integer, let $(\mathsf{Y}_{i},\mathcal{F}_{i})$ be a measurable space, and let $\psi_{i}:\mathsf{X}\to\mathsf{Y}_{i}$ be a measurable function. Let $M_{i}=\Pi\circ\psi_{i}^{-1}$, and let $(P_{i,y_{i}})_{y_{i}\in\mathsf{Y}_{i}}$ be a disintegration of $\Pi$ with respect to $\psi_{i}$. That is, (i) for $y_{i}\in\mathsf{Y}_{i}$, $P_{i,y_{i}}(\cdot)$ is a $\sigma$-finite measure on $(\mathsf{X},\mathcal{B})$, and, for $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$, $P_{i,y_{i}}(\cdot)$ is a probability measure concentrated on the pre-image $\psi^{-1}_{i}(y_{i}):=\psi^{-1}_{i}(\{y_{i}\})=\{x\in\mathsf{X}:\,\psi_{i}(x)=y_{i}\}$, (ii) for $f:\mathsf{X}\to[0,\infty)$ measurable, $y_{i}\mapsto\int_{\mathsf{X}}f(x)\,P_{i,y_{i}}(\mathrm{d}x)$ is measurable, and
@@
In practice, it may be difficult to sample exactly from $P_{i,y_{i}}(\cdot)$, preventing one from simulating the above Gibbs chain. Suppose that, instead, for $M_{i}$-a.e. $y_{i}\in\mathsf{Y}_{i}$ and $x\in\psi^{-1}_{i}(y_{i})$, it is possible to sample from $Q_{i,y_{i}}(x,\cdot)$, where $Q_{i,y_{i}}:\mathsf{X}\times\mathcal{B}\to[0,1]$ is an Mtk for $y_{i}\in\mathsf{Y}_{i}$, and it is reversible with respect to $P_{i,y_{i}}(\cdot)$ for $M_{i}$-a.e. $y_{i}$. This is the case for Metropolis-within-Gibbs algorithms, in which $Q_{i,y_{i}}$ is the (possibly multi-step) Mtk of some Metropolis-Hastings algorithm targeting $P_{i,y_{i}}$. Of course, if some $P_{i,y_{i}}(\cdot)$ is actually easy to sample from, one can simply take $Q_{i,y_{i}}(x,\cdot)=P_{i,y_{i}}(\cdot)$ for $x\in\mathsf{X}$. Throughout the text, we impose the following regularity condition:
@@ 2.2 A loose summary of main results
We now give an overview of our main results. The exact statements are given in Section 4, after a brief survey on the $L^{2}$ theory of Markov chains in Section 3. For now, recall that the $L^{2}$ norm of a Markov operator is in $[0,1]$. When the associated chain is reversible, the norm can be viewed as its convergence rate, with a smaller rate indicating faster convergence to the stationary distribution. It is also common to quantify the convergence speed of chains through the spectral gap, which is one minus the operator norm. The norms of $T$, $\hat{T}$, $S$, $\hat{S}$, and $Q_{i,y_{i}}$ are denoted by $\|T\|_{\Pi}$, $\|\hat{T}\|_{\Pi}$, $\|S\|_{M_{1}}$, $\|\hat{S}\|_{M_{1}}$, and $\|Q_{i,y_{i}}\|_{P_{i,y_{i}}}$, respectively.
For the random-scan Gibbs sampler, our bound loosely goes like this:
@@
$\left(\inf_{i}\text{Spectral gap of chain }C_{i}\right)\times\text{Spectral gap of chain }A\leq\text{Spectral gap of chain }B.$ (3)
Interestingly, bounds of this form have appeared in various other settings that are quite distinct from the one at hand (Carlen et al., 2003; Madras and Randall, 2002; Qin, 2023).
Previously, the only quantitative convergence bound for general hybrid random-scan Gibbs samplers that we know of is from Roberts and Rosenthal (1998). In the context of Example 3, Roberts and Rosenthal (1998) established a relation that bears some similarity to the first inequality in (2). However, their proof relied on the strong assumptions that the chain associated with $Q_{i,y_{i}}$ is uniformly ergodic for all $i$ and $y_{i}$, and the chain associated with $T$ is uniformly ergodic as well.
For hybrid data augmentation chains, a convergence bound of the form (3) was derived by Andrieu et al. (2018). Their bound essentially leads to the following set of inequalities:
$\left(1-\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\right)(1-\|S\|_{M_{1}})\leq 1-\|\hat{S}\|_{M_{1}}\leq\left(1+\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}\right)(1-\|S\|_{M_{1}}).$ (4)
If the Markov operator $Q_{2,y_{2}}$ is positive semi-definite for each $y_{2}$, then the right-hand-side can be tightened to $1-\|S\|_{M_{1}}$.
Remark 8.
Andrieu et al.’s (2018) bound is not identical to (4). The former involves an intractable quantity and concerns the right spectral gap instead of the absolute spectral gap. Bounding the intractable quantity in an obvious manner and slightly modifying the arguments in Andrieu et al. (2018) yields (4). The exact form of Andrieu et al.’s (2018) result will be given in Section 4.2.
For the lower bound on $1-\|\hat{S}\|_{M_{1}}$ in (4) to be non-trivial (strictly greater than 0), it is necessary that $\sup_{y_{2}}\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}<1$. Our contribution regarding hybrid data augmentation algorithms involves relaxing this requirement. By combining the arguments in Andrieu et al. (2018) and Łatuszyński and Rudolf (2014), we essentially derive that, for any positive even integer $t$,
$t(1-\|\hat{S}\|_{M_{1}})\geq 1-\|\hat{S}\|_{M_{1}}^{t}\geq 1-\|S\|_{M_{1}}-\sup_{y_{1}}\int_{\mathsf{X}}\|Q_{2,\psi_{2}(x)}\|_{P_{2,\psi_{2}(x)}}^{t}P_{1,y_{1}}(\mathrm{d}x).$ (5)
If $Q_{2,y_{2}}$ is positive semi-definite for each $y_{2}$, then (5) holds for odd values of $t$ as well. When $\|S\|_{M_{1}}<1$, the second inequality is nontrivial for large values of $t$ when
@@
That is, we no longer require $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$ to be bounded away from unity as $y_{2}$ varies; instead, we only need there to exist a set in $\mathcal{F}_{2}$ such that $\|Q_{2,y_{2}}\|_{P_{2,y_{2}}}$ is bounded away from 1 on that set, and the probability mass of that set evaluated by $P_{1,y_{1}}\circ\psi_{2}^{-1}(\cdot)$ is bounded away from 0 as $y_{1}$ varies.
@@
We were not able to obtain an analogue of (5) for the random-scan Gibbs chain that was sufficiently useful.
@@ 3 Preliminaries
This section reviews the $L^{2}$ theory for Markov chains.
Let $(\mathsf{Y},\mathcal{F},\omega)$ be a probability space. The linear space $L^{2}(\omega)$ consists of functions square-integrable function $f:\mathsf{Y}\to\mathbb{R}$ such that $\int_{\mathsf{Y}}f(y)^{2}\,\omega(\mathrm{d}y)<\infty$. Two functions in $L^{2}(\omega)$ are considered identical if they are equal almost everywhere. For $f,g\in L^{2}(\omega)$ and $c\in\mathbb{R}$, $(f+g)(y)=f(y)+g(y)$, and $(cf)(y)=cf(y)$. The inner product of $f$ and $g$ is $\langle f,g\rangle_{\omega}=\int_{\mathsf{Y}}f(y)g(y)\,\omega(\mathrm{d}y)$. The norm of $f$ is $\|f\|_{\omega}=\sqrt{\langle f,f\rangle_{\omega}}$. Let $L_{0}^{2}(\omega)$ be the subspace of $L^{2}(\omega)$ consisting of functions $f$ such that $\omega f:=\int_{\mathsf{Y}}f(y)\,\omega(\mathrm{d}y)=0$. It can be shown that $(L^{2}(\omega),\langle\cdot,\cdot,\rangle_{\omega})$ and $(L_{0}^{2}(\omega),\langle\cdot,\cdot,\rangle_{\omega})$ are real Hilbert spaces.
A probability measure $\mu:\mathcal{F}\to[0,1]$ is said to be in $L_{*}^{2}(\omega)$ if $\mathrm{d}\mu/\mathrm{d}\omega$ exists and is in $L^{2}(\omega)$. The $L^{2}$ distance between $\mu,\nu\in L_{*}^{2}(\omega)$ is
@@
where $\mu f$ and $\nu f$ are defined analogously to $\omega f$.
Let $K:\mathsf{Y}\times\mathcal{F}\to[0,1]$ be an Mtk, i.e., $K(x,\cdot)$ is a probability measure for $x\in\mathsf{X}$ and $K(\cdot,A)$ is a measurable function for $A\in\mathcal{F}$. Assume that
@@
We refer to $K$ as a Markov operator. The operator corresponding to the $t$-step Mtk $K^{t}(\cdot,\cdot)$ is precisely the operator $K^{t}=\prod_{i=1}^{t}K$. The operator $K$ is self-adjoint, i.e., $\langle Kf,g\rangle_{\omega}=\langle f,Kg\rangle_{\omega}$ for $f,g\in L_{0}^{2}(\omega)$, if and only if the corresponding chain is reversible with respect to $\omega$. We say $K$ is positive semi-definite if it is self-adjoint, and $\langle f,Kf\rangle_{\omega}\geq 0$ for $f\in L_{0}^{2}(\omega)$. With a bit of work, one can show that $T$ and $S$ from Section 2 are both positive semi-definite linear operators.
By Jensen’s inequality,